# Open-Source ESP32-C3 Wi-Fi Drivers for Static Analysis

### v2

@@ Acknowledgements.
- Schoeberl and Pedersen (2018) Martin Schoeberl and Rasmus Ulslev Pedersen. 2018. tpip: A time-predictable tcp/ip stack for cyber-physical systems. In 2018 IEEE 21st International Symposium on Real-Time Distributed Computing (ISORC). IEEE, IEEE Computer Society, 75–82. https://doi.org/10.1109/ISORC.2018.00018
- Schulz (2018) Matthias Schulz. 2018. Teaching Your Wireless Card New Tricks: Smartphone Performance and Security Enhancements Through Wi-Fi Firmware Modifications. Ph. D. Dissertation. Darmstadt University of Technology, Germany, Darmstadt. http://tuprints.ulb.tu-darmstadt.de/7243/
- Schuster et al. (2019) Simon Schuster, Peter Wägemann, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2019. Proving Real-Time Capability of Generic Operating Systems by System-Aware Timing Analysis. In Proceedings of the 25th Real-Time and Embedded Technology and Applications Symposium (RTAS ’19). IEEE, 318–330. https://doi.org/10.1109/RTAS.2019.00034
@@
- Systems (2016) Espressif Systems. 2016. ESP IoT Development Framework | Espressif Systems. https://www.espressif.com/en/products/sdks/esp-idf.
- Talla et al. (2017) Vamsi Talla, Bryce Kellogg, Shyamnath Gollakota, and Joshua R. Smith. 2017. Battery-Free Cellphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (Jun 2017), 25:1–25:20. https://doi.org/10.1145/3090090
- Vogelgesang et al. (2025) Kai Vogelgesang, Ishwar Mudraje, Luis Gerhorst, Phillip Raffeck, Peter Wägemann, Thorsten Herfet, and Wolfgang Schröder-Preikschat. 2025. PfIP: A UDP/IP Transactional Network Stack for Power-Failure Resilience in Embedded Systems. In Proceedings of the Consumer Communications & Networking Conference (CCNC 2025).
@@
- Yarahmadi and Rohou (2020) Bahram Yarahmadi and Erven Rohou. 2020. Compiler optimizations for safe insertion of checkpoints in intermittently powered systems. In Proceedings of the 20th International Conference on Embedded Computer Systems: Architectures, Modeling, and Simulation (SAMOS ’20), Vol. 12471. Springer, 169–185. https://doi.org/10.1007/978-3-030-60939-9_12

### v3

@@ Acknowledgements.
- Schoeberl and Pedersen (2018) Martin Schoeberl and Rasmus Ulslev Pedersen. 2018. tpip: A time-predictable tcp/ip stack for cyber-physical systems. In 2018 IEEE 21st International Symposium on Real-Time Distributed Computing (ISORC). IEEE, 75–82. https://doi.org/10.1109/ISORC.2018.00018
- Schulz (2018) Matthias Schulz. 2018. Teaching Your Wireless Card New Tricks: Smartphone Performance and Security Enhancements Through Wi-Fi Firmware Modifications. Ph. D. Dissertation. Darmstadt University of Technology, Germany. http://tuprints.ulb.tu-darmstadt.de/7243/
- Schuster et al. (2019) Simon Schuster, Peter Wägemann, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2019. Proving Real-Time Capability of Generic Operating Systems by System-Aware Timing Analysis. In Proceedings of the 25th Real-Time and Embedded Technology and Applications Symposium (RTAS ’19). 318–330. https://doi.org/10.1109/RTAS.2019.00034
@@
- Talla et al. (2017) Vamsi Talla, Bryce Kellogg, Shyamnath Gollakota, and Joshua R. Smith. 2017. Battery-Free Cellphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (2017), 25:1–25:20. https://doi.org/10.1145/3090090
- Vogelgesang et al. (2025) Kai Vogelgesang, Ishwar Mudraje, Luis Gerhorst, Phillip Raffeck, Peter Wägemann, Thorsten Herfet, and Wolfgang Schröder-Preikschat. 2025. PfIP: A UDP/IP Transactional Network Stack for Power-Failure Resilience in Embedded Systems. In Proceedings of the Consumer Communications & Networking Conference (CCNC 2025). https://www.nt.uni-saarland.de/publications/vogelgesang2025pfip/PfIP_authors_version.pdf
- Wägemann et al. (2018) Peter Wägemann, Christian Dietrich, Tobias Distler, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2018. Whole-System Worst-Case Energy-Consumption Analysis for Energy-Constrained Real-Time Systems. In Proceedings of the 30th Euromicro Conference on Real-Time Systems (ECRTS ’18), Vol. 106. 24:1–24:25. https://doi.org/10.4230/LIPIcs.ECRTS.2018.24
@@
- Yarahmadi and Rohou (2020) Bahram Yarahmadi and Erven Rohou. 2020. Compiler optimizations for safe insertion of checkpoints in intermittently powered systems. In Proceedings of the 20th International Conference on Embedded Computer Systems: Architectures, Modeling, and Simulation (SAMOS ’20), Vol. 12471. 169–185. https://doi.org/10.1007/978-3-030-60939-9_12